Fermi National Laboratory


Accelerator Update for June 17, 2005 - June 20, 2005

Friday June 17
The day shift began with the Antiproton Source (Pbar) stacking, with the Tevatron (TeV) in a dry squeeze, with the TeV separators in conditioning mode, and with MiniBooNE and NuMI taking beam.

Operations put the TeV into shot setup at 11:35 AM.

Operations established store 4214 at 1:56 PM with an initial luminosity of 68.19E30. Pbar resumed stacking and NuMI resumed taking beam.

Operations established beam to the Meson MCenter experiment at 8:07 PM.

Operations put the Recycler into shot setup at 8:55 PM to stash antiprotons. Pbar resumed stacking.

Saturday June 18
The midnight shift began with Operations monitoring store 4214 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:31 PM, two Main Injector power supplies tripped off (South and Center Anode). An operator reset the supplies, but then a RF station (MIRF9) tripped off and wouldn’t reset. With some expert advice, operators managed to get the station back up by 1:13 PM.

At 2:43 PM, Operations reported that MiniBooNE had been having trouble with their HVAC (air conditioning) all afternoon and asked to have their beam turned off due to equipment overheating. FESS people replaced a pump motor and MiniBooNE turned back on at 3:30 PM.

Operations terminated store 4214 at 4:19 PM.

Operations put the TeV into shot setup at 4:31 PM.

Operations established combination store 4215 at 6:30 PM with an initial luminosity of 107.3E30. NuMI resumed taking beam.

Operations resumed sending beam to MCenter at 6:42 PM.

At 6:50 PM, Operations reported that they could not resume stacking and called in a Pbar expert. The expert tuned the machine and established normal stacking by 8:53 PM.

Operations put the Recycler into shot setup to stash antiprotons at 11:04 PM.

At 11:08 PM, Operations received an alarm for the TeV sector F1 refrigeration building. Operators responded and stopped the wet engine. The TeV quenched at F1 at 11:13 PM. The store was lost. The stash was canceled.

MCenter resumed taking beam at 11:38 PM.

Pbar resumed stacking at 11:45 PM

Sunday June 19
The midnight shift began with the TeV recovering from a quench, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations began the TeV turn on at 6:25 AM.

Operations put the TeV through a dry squeeze at 7:07 AM.

Operations put the TeV into shot setup at 7:34 AM.

Operations established combination store 4217 at 9:22 AM with an initial luminosity of 82.4E30. Pbar resumed stacking and NuMI resumed taking beam.

MCenter resumed taking beam at 9:53 AM.

Recycler began stashing antiprotons at 1:28 PM. The stash ended about an hour later. Pbar resumed stacking.

MCenter resumed taking beam 2:41 PM.

Recycler began stashing antiprotons at 6:32 PM. The stash ended at 7:23 PM. Pbar resumed stacking and NuMI and MCenter resumed taking beam.

At 8:58 PM, MIRF5 tripped off for the first of seven times in one hour. It reset each time.

At 10:06 PM, a TeV sector B0 device failure (BB DI/DT) caused the loss of store 4217. There was no quench, the store aborted cleanly. Operators investigated and then called in an expert.

Monday June 20
The midnight shift began with TeV B0 work in progress, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The TeV expert had changed out the current regulator chassis and completed some cabling work on TeV sector B0 quadrupoles (B0Q2-4 and B0Q4) by 12:41 AM.

Operations put the TeV through a dry squeeze at 12:53 AM.

At 1:47 AM, the MI safety system tripped off on a ground fault indication. This trip caused the loss of the stash. Operations talked to an expert and then reset the trip at 1:54 AM.

Operations reestablished beam to Pbar, MiniBooNE, and NuMI at 2:03 AM.

Operations established beam to MCenter at 2:21 AM.

Operations put the TeV into shot setup at 3:36 AM.

Operations established store 4219 at 5:44 AM with an initial luminosity of 53.16E30.

Pbar resumed stacking at 6:25 AM.

MCenter resumed taking beam at 7 AM.

The Plan for Monday
The plans for today are to conduct Recycler stashing during the day and an end-of-store study this evening. TeV experts will then conduct another short study before the next store begins.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/20/2005   email Fermilab